OpenStudy (anonymous):

Given the function f(x) = 6(x+2) − 3, solve for the inverse function when x = 21.

OpenStudy (anonymous):

i have no idea what this means, but it could mean find \(f^{-1}(21)\)

OpenStudy (anonymous):

in which case, solve \[6(x+2)^2-3=21\] for \(x\)

OpenStudy (anonymous):

ohk

hero (hero):

@preddie1234, have you found \(f^{-1}(21)\) yet?

OpenStudy (anonymous):

no im trying to

hero (hero):

What have you done so far?

OpenStudy (anonymous):

add the things together

hero (hero):

Add what together?

OpenStudy (anonymous):

the equation

hero (hero):

How does doing that help you find the inverse? Show the work you have done so far.

OpenStudy (anonymous):

it dont look right

hero (hero):

So you have f(x) = 6(x+2) − 3 Change f(x) to y first.

OpenStudy (anonymous):

oh k so that will be y=6x+12-3

hero (hero):

Now swap x and y

OpenStudy (anonymous):

oh what next

hero (hero):

Did you swap x and y already?

OpenStudy (anonymous):

yes

geerky42 (geerky42):

How did you get 6(x+2)^2−3=21? @satellite73

hero (hero):

@preddie1234, what did you get after swapping x and y?

hero (hero):

In case you haven't noticed, I'm interested in seeing the work you do as you are doing it.

OpenStudy (shamim):

plz see my attachment

hero (hero):

@shamim, no offense, but that's barely readable. Furthermore, you should be trying to guide the user rather than post everything all at once.